OpenStudy (anonymous):

One circular table has a diameter of 9 ft, and another circular table has a diameter of 20 ft. How much greater is the area of the larger table? Round to the nearest whole number

OpenStudy (anonymous):

\[\frac{\pi r^2\text{/.}\, r\to 20}{\pi r^2\text{/.}\, r\to 9}=\frac{400}{81}=4.93827=5 \] The above is a Mathematica expression. What it says is that if r in the numberator is replaced by 20 the the same sort of thing for the denominator, the ratio of the areas is 400/81 or 5 rounded.

OpenStudy (anonymous):

numberator should have been spelled, numerator. Sorry.